DIRC Operations Report

Period from 2008/02/19 to 2008/03/19


Oncall: Jennifer Prendki, Jose Benitez
Fast Monitoring: Jennifer Prendki, Jose Benitez

Contents:

    PEP     BaBar     DIRC     Backgrounds     Scheduled activities     Pending/Ongoing Work     Links

Luminosity:

PEP peak lumi: ~12.0 x 1033

Run 7 total delivery: ~45.90 fb-1

with : ~0.62 fb-1 on the Y(4S), ~30.66 fb-1 on the Y(3S), ~12.12 fb-1 on the Y(2S).

PEP, BaBar activities/issues since last DIRC meeting

  • Took some Y(3S) offpeak data.
  • Switched to running on the Y(2S) on March 1st.
  • Backgrounds are high since start of Y(2S) data.
  • Took some Y(2S) offpeak data on Tuesday.

    DIRC activities/issues since last DIRC meeting

  • Shoulder in timing distributions was back on Monday Swing and Tuesday Day shifts. Went worse when switching to Y(2S) offpeak. Fast monitoring plots showed trigger lines #1 & #8 were responsible for that.
  • Current detector status :
  • One chip in Sector #0 seems to be dying those days. We keep an eye on this.

    See DIRC logbook for details

    Lumi and Background History

  • Some more plots for background history :

    Spare status

    Equipment # used Spares
    DFB 168 12/17 good + 4 possibly good, 1 bad
    DCC 12 4/5 good
    HV Boards 48 9/11 good , 2 broken recovered by Jose from Angels shop because they couldn't fix them

    Manpower situation

    Links